# DUNIA: Pixel-Sized Embeddings via Cross-Modal Alignment for Earth Observation Applications

### v1

@@ E.2 Using a Shared Decoder
E.3 Hierarchical VICReg Loss
The original VICReg loss proposed by Bardes et al. 2021 was applied between two views of the same input image. They later extended their work and applied it to local image features (Bardes et al., 2022). Their results showed that better results were obtained when applying the VICReg loss on local features and at the instance level. Here, we further extend this work and apply it hierarchically to the decoder outputs between the pre-training model and the multi-temporal AE. Results in Table 5 show significant improvements for both datasets when using this formulation. These results are also in line with the findings in Bardes et al. 2022.
Table 5: Zero-shot classification performance ($wF1$) on the $PF$ and $CLC_{+}$ datasets using a model trained with multiple VICReg losses at different decoder levels ($\text{VICReg}_{h}$) against a single loss applied on the pixel-level embeddings from the last decoder layer ($\text{VICReg}_{s}$).
@@ E.3 Hierarchical VICReg Loss
$\text{VICReg}_{h}$ 76.0% 80.1%
$\text{VICReg}_{s}$ 67.6% 74.2%
E.4 Neighborhood Attention
The output of our pre-training model is composed of two neighborhood attention layers per output head, instead of the standard convolutional layers. This design choice allows each embedding to be modeled based on its local neighborhood instead of relying on small and fixed receptive fields. Table 6 shows an increase of 2.1% ($wF1$) for the $CLC_{+}$ dataset and an $rmse$ decrease of 0.4 m for the $\mathcal{W}_{rh}$ dataset using this new configuration.
Table 6: Performance differences between a convolutional output layer (CNN) vs. a NA layer on the $CLC_{+}$ and $\mathcal{W}_{rh}$ datasets.
@@ E.4 Neighborhood Attention
CNN 72.1% ($wF1$) 2.4 ($rmse$)%
NA 74.2% ($wF1$) 2.0 ($rmse$)%
E.5 Embedding sensitivity to horizontal and vertical structures
Our results, either in the zero-shot setting or the fine-tuning setting used embeddings from $\mathcal{O}^{\mathcal{H}}$ for horizontal structure-related products (e.g., tree species identification, land cover mapping) and embeddings from $\mathcal{O}^{\mathcal{V}}$ for vertical structure-related products (e.g., canopy height mapping, waveform retrieval). We perform the same tests in the zero-shot setting, but reversing the query and retrieval embeddings for the other structure direction. Table 7 shows that relying on vertical structure data for products requiring horizontal understanding and vice versa performs poorly for the six tested products. This validates our design choice of cross-modal alignment with vertical structure data for Earth Observation products relying on this type of information.
Table 7: Top-1 retrieval-based zero-shot classification performance of DUNIA with opposite direction query and retrieval embeddings. OG represents the original results from Table 1 for S samples and KNN=5. and are DUNIA query embeddings from $\mathcal{O^{V}}$ and $\mathcal{O^{H}}$ respectively.
@@ E.5 Embedding sensitivity to horizontal and vertical structures
$CLC_{+}$ $wF1$ 80.1% 35.1%
$PASTIS$ $OA$ 56.2% 0.0%
$PF$ $wF1$ 73.8% 56.1%
Appendix F Further Results
@@
Figure 3: Maps produced through different models for three products: fractional canopy cover ($\mathcal{W}_{c}$), canopy height ($\mathcal{W}_{r}$), and land cover classes ($CLC_{+}$). Baseline represents reference maps from respectively: adapted FORMS (Schwartz et al., 2023) for $\mathcal{W}_{c}$, FORMS (Schwartz et al., 2023) for $\mathcal{W}_{rh}$ and $CLC_{+}$ (European Environment Agency, 2024). DUNIA (ZS) represents maps obtained in the zero-shot setting, while DUNIA (FT) represents maps obtained in the fined-tuned setting. Best viewed zoomed-in (300+%).
@@
Figure 4: Comparison of map qualities for the task of estimating the canopy height in the zero-shot setting using 50K $l$ GEDI samples as queries (left) vs. only 10K $l$ (right).
@@
Figure 5: Uncurated list of retrieved ( ) and generated waveforms ( ) overlayed on a reference waveform ( ). ’Correlation’ is Pearson’s correlation coefficient (r) between the reference waveform and the retrieved/generated waveform. Best viewed zoomed-in (200+%)
@@
Figure 6: Uncurated list of retrieved ( ) and generated waveforms ( ) overlayed on a reference waveform ( ). ’Correlation’ is Pearson’s correlation coefficient (r) between the reference waveform and the retrieved/generated waveform. Best viewed zoomed-in (200+%)
@@
Figure 7: Uncurated list of retrieved ( ) and generated waveforms ( ) overlayed on a reference waveform ( ). ’Correlation’ is Pearson’s correlation coefficient (r) between the reference waveform and the retrieved/generated waveform. Best viewed zoomed-in (200+%)
@@
Figure 8: Uncurated list of retrieved ( ) and generated waveforms ( ) overlayed on a reference waveform ( ). ’Correlation’ is Pearson’s correlation coefficient (r) between the reference waveform and the retrieved/generated waveform. Best viewed zoomed-in (200+%)

### v2

@@ H.2 Using a Shared Decoder
H.3 Hierarchical VICReg Loss
The original VICReg loss proposed by Bardes et al. (2021) was applied between two views of the same input image. They later extended their work and applied it to local image features (Bardes et al., 2022). Their results showed that better results were obtained when applying the VICReg loss on local features and at the instance level. Here, we further extend this work and apply it hierarchically to the decoder outputs between the pre-trained model and the multi-temporal AE. Results in Table H3 show significant improvements for both datasets when using this formulation. These results are also in line with the findings in Bardes et al. (2022).
Table H3: Zero-shot classification performance ($w$F1) on the $PF$ and $CLC_{+}$ datasets using a model pre-trained with multiple VICReg losses at different decoder levels ($\text{VICReg}_{h}$) against a single loss applied on the embeddings from the last decoder layer ($\text{VICReg}_{s}$).
@@ H.3 Hierarchical VICReg Loss
$\text{VICReg}_{h}$ 76.0 80.1
$\text{VICReg}_{s}$ 67.6 74.2
H.4 Neighborhood Attention
The output of our pre-trained model is composed of two neighborhood attention layers per output head, instead of the standard convolutional layers. This design choice allows each embedding to be modeled based on its local neighborhood instead of relying on small and fixed receptive fields. Table H4 shows an increase of 2.1% ($w$F1) for the $CLC_{+}$ dataset and an RMSE decrease of 0.4 m for the $\mathcal{W}_{rh}$ dataset using this new configuration.
Table H4: Performance differences between a convolutional output layer (CNN) vs. a NA layer on the $CLC_{+}$ and $\mathcal{W}_{rh}$ datasets.
@@ H.4 Neighborhood Attention
CNN 72.1 (wF1) 2.4 (RMSE)
NA 74.2 (wF1) 2.0 (RMSE)
H.5 Embedding sensitivity
Our results, either in the zero-shot setting or the fine-tuned setting used embeddings from $\mathcal{O}^{\mathcal{H}}$ for horizontal structure-related products (e.g., tree species identification, land cover mapping) and embeddings from $\mathcal{O}^{\mathcal{V}}$ for vertical structure-related products (e.g., canopy height mapping, waveform retrieval). We perform the same tests in the zero-shot setting, but reversing the query and retrieval embeddings for the other structure direction. Table H5 shows that relying on vertical structure data for products requiring horizontal understanding and vice versa performs poorly for the six tested products. This validates our design choice of cross-modal alignment with vertical structure data for EO products relying on this type of information.
Table H5: Top-1 retrieval-based zero-shot classification performance of DUNIA with opposite direction query and retrieval embeddings. OG represents the original results from Table 1 for KNN=50. and are DUNIA query embeddings from $\mathcal{O^{V}}$ and $\mathcal{O^{H}}$ respectively.
@@ H.5 Embedding sensitivity
$CLC_{+}$ wF1 80.1 35.1
$PASTIS$ OA 56.2 0.0
$PF$ wF1 73.8 56.1
H.6 Inference Runtime (Zero-Shot)
Generating maps in the zero-shot setting requires: 1) a forward pass through the pre-trained model (encoder-dual decoders) to obtain the embeddings, 2) for each pixel, retrieve the $k$ nearest neighbors (KNN) from the retrieval database (DB), and 3) classify the queried pixel embedding based on distance-weighted voting. As such, generating a map over e.g. a $20.48\times 20.48$ Km area at 10 m resolution, requires querying and classifying 4,194,304 pixels. However, Table H6 shows that the querying and classification times only add a small overhead. Moreover, the time required for the three operations (forward pass, retrieval, and classification) is still faster than a similarly capable model like AnySat.
Table H6: Inference runtime (in seconds) of DUNIA in the zero-shot setting, compared to AnySat. DB denotes the retrieval database of $N$ key/value pairs (embedding/label), and KNN is the number of nearest neighbors retrieved per pixel. All times are in seconds.
Model DB KNN Forward pass Retrieval Classification Total
DUNIA 256K 100 2.52 0.36 1.34 4.22
DUNIA 512K 200 2.52 0.40 1.88 4.80
AnySat 177.37 177.37
H.7 Impact of Input Size
To assess the sensitivity of DUNIA to the input image resolution, we conducted experiments using three different input sizes: $128\times 128$, $256\times 256$, and $512\times 512$. As shown in Table H7, the results across datasets and metrics remain unchanged. This confirms that image size has a negligible impact on the final performance of the model in the zero-shot setting.
Table H7: Zero-shot performance of DUNIA at different input image sizes for several datasets.
Dataset Metric 128$\times$128 256$\times$256 512$\times$512
$\mathcal{W}_{rh}$ RMSE (r) 2.2 2.0 2.1
$\mathcal{W}_{c}$ RMSE (r) 11.6 11.7 11.7
$CLC+$ wF1 80.2 80.1 80.2
H.8 Temporal Stability
To evaluate the temporal stability of DUNIA, we conducted two experiments on several vertical structural products across the years 2019, 2020, and 2021. Other products were excluded due to the unavailability of labels for different years.
1. Fine-tuning an MLP head on top of a model pre-trained in 2020, and evaluating on 2019 and 2021.
2. Pre-training DUNIA using combined data from 2019–2021 and evaluating it in the zero-shot setting.
Table H8 shows that fine-tuning on individual years yields consistent performance across time. Similarly, Table H9 demonstrates that zero-shot performance is also stable over several years.
Table H8: Fine-tuned performance (RMSE) on vertical structure variables estimation across several years.
Dataset Metric 2019 2020 2021
$\mathcal{W}_{rh}$ RMSE 1.35 1.34 1.40
$\mathcal{W}_{c}$ RMSE 9.8 9.8 10.1
$\mathcal{W}_{pai}$ RMSE 0.65 0.62 0.63
Table H9: Zero-shot performance (RMSE) on vertical structure variables retrieval across years.
Dataset Metric 2019 2020 2021
$\mathcal{W}_{rh}$ RMSE 2.4 2.0 2.1
$\mathcal{W}_{c}$ RMSE 11.6 11.7 11.9
$\mathcal{W}_{pai}$ RMSE 0.75 0.71 0.74
H.9 Impact of S-2 and S-2 Input Images
Table H10 shows that combining both modalities yields the best performance for height estimation and tree species identification, while land cover classification is relatively robust to the choice of modality.
Table H10: Performance of DUNIA when using S-1 only, S-2 only, or both modalities. Results are reported for height prediction (RMSE), land cover classification ($w$F1) on the $CLC_{+}$ dataset, and tree species identification ($w$F1) on the $PF$ dataset. Performance results obtained in the fine-tuned setting.
Dataset Metric S-1 Only S-2 Only S-2 & 1
$\mathcal{W}_{rh}$ RMSE 3.80 2.80 1.34
$PF$ wF1 65.5% 81.2% 82.2%
$CLC_{+}$ wF1 90.2% 90.0% 90.3%
H.10 Median Composites vs. Single-Date Images
Results on height estimation and crop classification ($PASTIS$) (Table H11) show that median composites — even without full time series — significantly outperform single-date imagery by capturing more stable and representative reflectance patterns over time.
Table H11: Comparison between single-date and median composite inputs for products where temporal information is critical. Performance results obtained in the fine-tuned setting.
Dataset Metric Single-date image Median composite
$\mathcal{W}_{rh}$ RMSE 1.9 1.3
$PASTIS$ wF1 42.3 77.0
Appendix I Further Results
@@
Figure I1: Fractional canopy cover ($\mathcal{W}_{c}$) maps produced with different models. Baseline represents reference maps from adapted FORMS (Schwartz et al., 2023). DUNIA (ZS) represents maps obtained in the zero-shot setting, while DUNIA (FT) represents maps obtained in the fined-tuned setting. Best viewed zoomed-in (300+%).
Figure I2: Canopy height ($\mathcal{W}_{rh}$) maps produced with different models. Baseline represents reference maps from FORMS (Schwartz et al., 2023). DUNIA (ZS) represents maps obtained in the zero-shot setting, while DUNIA (FT) represents maps obtained in the fined-tuned setting. Best viewed zoomed-in (300+%).
Figure I3: Land cover classes ($CLC_{+}$) maps produced with different models. Baseline represents reference maps from the European Environment Agency (2024). DUNIA (ZS) represents maps obtained in the zero-shot setting, while DUNIA (FT) represents maps obtained in the fined-tuned setting. Best viewed zoomed-in (300+%).
@@
Figure I4: Comparison of map qualities for the task of estimating the canopy height in the zero-shot setting using 50K $l$ GEDI samples as queries (top row) vs. only 10K $l$ (bottom row).
@@
Figure I5: Uncurated list of retrieved ( ) and generated waveforms ( ) overlayed on a reference waveform ( ). ’Correlation’ is Pearson’s correlation coefficient (r) between the reference and the retrieved/generated waveforms. Best viewed zoomed-in (200+%).
@@
Figure I6: Uncurated list of retrieved ( ) and generated waveforms ( ) overlayed on a reference waveform ( ). ’Correlation’ is Pearson’s correlation coefficient (r) between the reference and the retrieved/generated waveforms. Best viewed zoomed-in (200+%).
@@
Figure I7: Uncurated list of retrieved ( ) and generated waveforms ( ) overlayed on a reference waveform ( ). ’Correlation’ is Pearson’s correlation coefficient (r) between the reference and the retrieved/generated waveforms. Best viewed zoomed-in (200+%).
@@
Figure I8: Uncurated list of retrieved ( ) and generated waveforms ( ) overlayed on a reference waveform ( ). ’Correlation’ is Pearson’s correlation coefficient (r) between the reference and the retrieved/generated waveforms. Best viewed zoomed-in (200+%).